# Asymmetric coupling of non-chaotic Rulkov neurons: fractal attractors, quasi-multistability, and final state sensitivity

### v1

@@ III Geometry of the chaotic pseudo-attractor
We now present the results from our analysis of the non-chaotic spiking attractor and the chaotic spiking-bursting pseudo-attractor’s geometry. In Figs. 3 and 4, we present visualizations of these two attractors by plotting a subset of the orbit of the initial states $\mathbf{X}_{0}=(-0.54,-3.25,-1,-3.25)$ for the non-chaotic attractor and $\mathbf{X}_{0}=(-0.56,-3.25,-1,-3.25)$ for the chaotic pseudo-attractor. In order to visualize these attractors embedded in four-dimensional space $(x_{1},y_{1},x_{2},y_{2})$, we plot their projections onto the $(x_{1},y_{1})$ and $(x_{2},y_{2})$ planes. It is immediately clear that the spiking attractor (Fig. 3) is associated with a non-chaotic periodic orbit and the pseudo-attractor (Fig. 4) is associated with complex chaotic orbits.
Due to the apparent geometrical complexity of the chaotic pseudo-attractor in Fig. 4, we suspect that the pseudo-attractor is fractal and strange. However, the fact that orbits on the pseudo-attractor are chaotic does not guarentee that the pseudo-attractor is fractal, as there exist chaotic attractors that are not strange [53]. The non-chaotic Rulkov map lends itself especially to the possiblity of a chaotic attractor that isn’t strange due to its resetting mechanism that maps all sufficiently high $x$ values back to $-1$ (Eq. 2), which always results in at least one Lyapunov exponent of $\infty$- ∞ [49]. Therefore, we perform a direct computation of the fractal dimension of the chaotic pseudo-attractor using the standard box-counting algorithm [54]. Using four-dimensional boxes of size $\epsilon=1/20,1/30,1/40$, we get that the box-counting dimension of the chaotic pseudo-attractor is $d\approx 1.84$ with an $R^{2}$ value of 0.9999, indicating that the chaotic pseudo-attractor is indeed fractal and strange. However, the pseudo-attractor’s dimension is still quite small for a geometric object embedded in four-dimensional space, which we attribute to the non-chaotic silence between bursts of spikes (see Fig. 1(b) and the left sides of Figs. 4(a) and 4(b)), as well as the aforementioned resetting mechanism.
We also calculate the Lyapunov dimension
@@
and we assume the Lyapunov exponents are ordered from greatest to least [55]. The Kaplan-Yorke conjecture states that the Lyapunov dimension of typical attractors is equal to their fractal dimension [56]. For the chaotic pseudo-attractor orbit displayed in Fig. 4, we compute its Lyapunov spectrum to be approximately $\lambda\approx\{0.0165,-0.0105,-0.0810,-\infty\}$. Then, the Lyapunov dimension of the chaotic pseudo-attractor is approximately $d_{l}\approx 2.07$, which is reasonably far from the pseudo-attractor’s true fractal dimension. We believe that this is the case because these chaotic dynamics are occurring on a pseudo-attractor and not a true attractor, for which the Kaplan-Yorke conjecture almost always holds [57].
(a) Maximal Lyapunov exponent color map
(b) Chaotic (white) and non-chaotic (black) basins
@@ IV Basin classification
We have established that a quasi-multistability exists in this asymmetrically coupled neuron system, so we are now interested in analyzing the geometrical properties of the basins of attraction of the non-chaotic spiking attractor and strange pseudo-attractor. We begin by classifying the basins of attraction using the method of Sprott and Xiong [27], which we briefly outline here. Let $A=\{\mathbf{a}_{1},\mathbf{a}_{2},\ldots\}$ be an attractor  [^2] We assume $A$ is composed of an infinite number of points, but this method can be trivially altered for attractors of a finite number of points.. Define the mean or “center of mass” of $A$ to be
@@ IV.1 Two-dimensional basin slices
We now need a way to determine which attractor an initial state $\mathbf{X}_{0}\in S_{2}^{\prime}$ is initially attracted to. This is nontrivial because the non-chaotic spiking attractor and chaotic pseudo-attractor occupy similar regions of state space (see Figs. 3 and 4). Fig. 2 illustrates that states attracted to the chaotic pseudo-attractor spend on the order of $10^{4}$ iterations in the pseudo-attractor before going to the spiking attractor. We find that testing other initial states yields similar results. Therefore, we use an indirect method of determining which basin an initial state is in, saying that an initial state that has a positive maximal Lyapunov exponent after 5,000 iterations is attracted to the chaotic pseudo-attractor, while an initial state that has a negative $\lambda_{1}$ after the same amount of time is attracted to the spiking attractor.
Using this method on a large number of initial states $\mathbf{X}_{0}\in S_{2}^{\prime}$, we present our maximal Lyapunov exponent results with a color map in Fig. 5(a). An immediate observation is the beautiful complexity of the different Lyapunov exponents in state space. This complexity is further emphasized when we color the basins properly in Fig. 5(b), with the basin of the chaotic pseudo-attractor (positive $\lambda_{1}$) shown in white and the basin of the non-chaotic attractor (negative $\lambda_{1}$) shown in black. Here, we can see that most of $S_{2}^{\prime}$ is taken up by the basin of the white chaotic pseudo-attractor, with some black regions, curves, and scattered points of stability scattered throughout. The most obvious of these features is the black line that goes across the diagonal of Fig. 5(b) from the bottom left to the top right. This line and its surrounding area are associated with initial states where the two individual neurons’ states are the same or very close to being the same. In this case, it makes sense that the neurons will immediately be synchronized, spiking together in the non-chaotic attractor.
@@
Because $S_{2}$ is two-dimensional, we also need to specify what is meant by the mean and standard deviation of the four-dimensional attractors within the two-dimensional plane $S_{2}$. Let us denote a generic four-dimensional attractor as $A_{4}$ and its associated basin as $\hat{A}_{4}$  [^3] In the asymmetrically coupled Rulkov neuron system, $A_{4}$ could denote either the non-chaotic spiking attractor or the chaotic pseudo-attractor.. From Eq. 14, if
@@
which is the projection of $\langle A_{4}\rangle$ onto the $(x_{1},x_{2})$ plane. Similarly, the true standard deviation of $A_{4}$ in four-dimensional state space is, from Equation 15,
@@
Following Eq. 16, for our classification of the basin slices $\hat{A}_{4}\cap S_{2}$, we will consider the normalized two-dimensional distance $\xi_{2}$ from $\langle A_{2}\rangle$ of some $\mathbf{X}=(x_{1},-3.25,x_{2},-3.25)\in S_{2}$ to be,
@@
To classify the basin slices of the chaotic pseudo-attractor and non-chaotic spiking attractor, we want to find the two probability functions $P_{w}(\xi_{2})$ and $P_{b}(\xi_{2})$ associated with the white (chaotic) and black (non-chaotic) basins, respectively, which we accomplish using an iterative Monte Carlo method. Specifically, we first randomly select states from the open disk of radius $\xi_{2}=1$ centered at $\langle A_{2}\rangle$ and calculate the maximal Lyapunov exponents of the initial states’ orbits to determine which basin they end up in. Using this, we calculate the value of $P(1)$. Then, starting from $k=0$, we repeat the same process for the two-dimensional shell centered at $\langle A_{2}\rangle$ with inner radius $\xi_{2}=2^{k}$ and outer radius $\xi_{2}=2^{k+1}$ to calculate the value of $\Delta P(2^{k})$. Then, we use Eq. 21 to find $P(2^{k+1})$ and repeat the process for $k+1$. One important numerical note is that, given we are testing values of $\xi_{2}=2^{k}$, the maximum value of $k$ must be chosen with care. On the one hand, since we are interested in the limit $\xi\to\infty$, $k$ must be large enough that we capture the geometry of the basins far away from the attractors. On the other hand, since we are only iterating 5000 times, it also must be small enough that initial states go to an attractor fast enough to get an accurate Lyapunov exponent calculation on it. By computer experiment, we determine that a good maximum value of $k$ is 11.
@@
Table 1: Some approximate $P(\xi_{2})$ values of the asymmetrically coupled Rulkov neuron system’s chaotic (white) and non-chaotic (black) basins
In Table 1, we display our numerical results for the values of $P_{w}(\xi_{2})$ and $P_{b}(\xi_{2})$, where we neglect values of $k$ from 0 to 5 since we are interested in large values of $\xi_{2}$. We can immediately see that $P_{w}(\xi_{2})$ is behaving a bit oddly, as Equation 20 indicates that $P(\xi)$ functions should either decrease or stay the same. For this reason, we first turn our attention to $P_{b}(\xi_{2})$, which is behaving as expected. Performing a linear regression calculation on $\log_{2}P_{b}(\xi_{2})$ against $\log_{2}\xi_{2}$, we get that
@@
Since $\gamma\approx 0.2957$ is between 0 and $n=2$, we can conclude that the basin slice of the non-chaotic spiking attractor is Class 3, meaning it extends to infinity in some directions but takes up increasingly small fractions of state space as we move outwards. This makes sense because the farther away an initial state is from the non-chaotic spiking attractor, the lower the chance there will be for the neurons to immediately synchronize with each other. In the limit $\xi_{2}\to\infty$, we expect that the only initial conditions that get attracted to the spiking attractor immediately are the ones where $x_{1,0}=x_{2,0}$, which is the diagonal in Figure 5. For this reason, we can conclude that the black basin slice in $S_{2}$ has finite measure since the line $x_{1,0}=x_{2,0}$ doesn’t contribute to the measure of the two-dimensional black basin slice.
@@
by the definitions of the spiking attractor and chaotic pseudo-attractor: if an initial state does not immediately fall into a non-chaotic spiking orbit with negative $\lambda_{1}$, then it goes to a chaotic orbit with positive $\lambda_{1}$. This fact can be confirmed by adding the data horizontally across Table 1, which gives values very close to 1. Therefore, even though $P_{w}(2^{11})$ is less than $P_{w}(2^{10})$ due to the aforementioned difficulty of calculating Lyapunov exponents for orbits of initial states far away from their attractors, the values of $P_{w}(\xi_{2})$ are indeed approaching 1. This can be seen by exponentiating both sides of Equation 30:
@@ IV.2 Four-dimensional basins
using the definitions from Eqs. 25 and 27. This is similar to the process for the two-dimensional slices, with a few noteworthy changes. One nontrivial aspect of this is choosing a random state $\mathbf{X}_{0}$ from a four-dimensional open ball with radius $\xi_{4}$ centered at $\langle A_{4}\rangle$. In two-dimensional space with an open disk $|\mathbf{x}-\langle A_{2}\rangle|<\xi_{2}$, we simply choose a random $r\in[0,\xi_{2})$ and a random $\phi\in[0,2\pi)$, then say the random initial state is $\mathbf{x}_{0}=\langle A_{2}\rangle+(r\cos\phi,r\sin\phi)$. In four-dimensional space, we can use the analog to spherical coordinates in four-dimensions, where four coordinates $r\in[0,\xi_{4})$, $\theta_{1}\in[0,\pi)$, $\theta_{2}\in[0,\pi)$, and $\phi\in[0,2\pi)$ are chosen randomly. Then, a random initial state $\mathbf{X}_{0}\in S(\xi_{4})$ is given by [60]
@@
Table 2: Some approximate $P(\xi_{4})$ values of asymmetrically coupled Rulkov neuron system’s chaotic (white) and non-chaotic (black) basins
@@
(b) Chaotic (white) and non-chaotic (black) basins
@@
The results for the probability function values of the white basin $P_{w}(\xi_{4})$ are shown in Table 2. In the table, the $P_{b}(\xi_{4})$ values are indirectly calculated using $P_{b}(\xi_{4})=1-P_{w}(\xi_{4})$, which follows from Equation 31. Observing the values in Table 2, it appears that $P_{w}(\xi_{4})$ and $P_{b}(\xi_{4})$ remain relatively constant as $\xi_{4}$ is varied, which is a rather unexpected result given our analysis of the two-dimensional basin slices. Running linear regressions on the log-log plots of both $P_{w}(\xi_{4})$ and $P_{b}(\xi_{4})$, we find that
@@
with $R^{2}=0.386$. These low $R^{2}$ values confirm that there is no upward or downward trend in $P_{w}(\xi_{4})$ or $P_{b}(\xi_{4})$. This means that, within the numerical variability of the Monte Carlo method, the value of the basin classification parameter $\gamma$ is effectively 0 for both the white and black basins. From Eqs. 35 and 36, the value of the other basin classification parameter $P_{0}$ is $P_{0}=0.8517$ for the white basin and $P_{0}=0.1482$ for the black basin. Both of these are between 0 and 1, so both the white and black basins are Class 2, meaning they both occupy a fixed fraction of four-dimensional state space. This makes sense considering the data in Table 2, but again, it is an unexpected result.
@@
Our explanation for why the white and black basins are Class 2 relies on the fact that the non-chaotic Rulkov map is a slow-fast system. Because $x$ is a fast variable and $y$ is a slow variable, $x_{1}$ and $x_{2}$ adjust much faster than $y_{1}$ and $y_{2}$, so if we consider some initial state $\mathbf{X}_{0}=(x_{1,0},y_{1,0},x_{2,0},y_{2,0})$ some distance $\xi_{4}$ away from either attractor $A_{4}$, $y_{1}$ and $y_{2}$ will slowly drift towards $\langle y_{1}\rangle$ and $\langle y_{2}\rangle$, respectively, and $x_{1}$ and $x_{2}$ will easily keep up with their slow attraction. By the time $y_{1}$ and $y_{2}$ reach the region around $\langle y_{1}\rangle$ and $\langle y_{2}\rangle$, $x_{1}$ and $x_{2}$ will also be near $\langle x_{1}\rangle$ and $\langle x_{2}\rangle$, respectively. This will occur regardless of how much farther away $x_{1,0}$ and $x_{2,0}$ are from $\langle x_{1}\rangle$ and $\langle x_{2}\rangle$ compared to the distance from $y_{1,0}$ and $y_{2,0}$ to $\langle y_{1}\rangle$ and $\langle y_{2}\rangle$, respectively, since the fast variables $x_{1}$ and $x_{2}$ evolve so much more quickly than the slow variables $y_{1}$ and $y_{2}$. Thus, once $y_{1}$ and $y_{2}$ get close to the attractors, we are effectively starting in $S_{2}^{\prime}$ (or a square of similar size close and parallel/skew to it). Indeed, comparing the fractions of state space that the white ($P_{0}=0.8517$) and black ($P_{0}=0.1482$) basins take up to the values of $P_{w}(\xi_{2})$ and $P_{b}(\xi_{2})$ for the small values of $\xi_{2}$ in Table 1, we find that these align quite well.
@@
Although we cannot visualize the entirety of these four-dimensional basins, we can get a grasp of this Class 2 basin behavior by graphing a different two-dimensional slice of the basins, namely, one where $y_{1,0}$ and $y_{2,0}$ are varied but $x_{1,0}$ and $x_{2,0}$ are fixed. In Fig. 6, we graph a large range of initial $y$ values $-256\leq y_{1,0}\leq 256$ and $-256\leq y_{2,0}\leq 256$ while fixing the initial $x$ values at $x_{1,0}=-1$ and $x_{2,0}=1$. In Fig. 6(b), we can see a dispersed, seemingly random distribution of white and black basin points with no defined pattern in their arrangement. Similarly, in Fig. 6(a), the maximal Lyapunov exponent also appears random. The important observation from Fig. 6(b) is that the distribution of white and black points appears to remain the same everywhere on the graph, despite the fact that it spans a large range of different $y$ values. This is characteristic of Class 2 basins: the fraction of state space the basins take up doesn’t depend on the distance from their attractors. Fig. 6(b) also lacks the black diagonal that appears in Figure 5(b). This line of stability doesn’t appear in this graph because, as previously mentioned, it is defined by the equation

### v3

@@ III Geometry of the chaotic pseudo-attractor
We now present the results from our analysis of the nonchaotic spiking attractor and the chaotic spiking-bursting pseudo-attractor’s geometry. In Figs. 3 and 4, we present visualizations of these two attractors by plotting a subset of the orbit of the initial states $\mathbf{X}_{0}=(-0.54,-3.25,-1,-3.25)$ for the nonchaotic attractor and $\mathbf{X}_{0}=(-0.56,-3.25,-1,-3.25)$ for the chaotic pseudo-attractor. In order to visualize these attractors embedded in four-dimensional space $(x_{1},y_{1},x_{2},y_{2})$, we plot their projections onto the $(x_{1},y_{1})$ and $(x_{2},y_{2})$ planes. It is immediately clear that the spiking attractor (Fig. 3) is associated with a nonchaotic periodic orbit and the pseudo-attractor (Fig. 4) is associated with complex chaotic orbits.
Due to the apparent geometrical complexity of the chaotic pseudo-attractor in Fig. 4, we suspect that the pseudo-attractor is fractal and strange. However, the fact that orbits on the pseudo-attractor are chaotic does not guarantee that the pseudo-attractor is fractal, as there exist chaotic attractors that are not strange [53]. The nonchaotic Rulkov map lends itself especially to the possibility of a chaotic attractor that isn’t strange due to its resetting mechanism that maps all sufficiently high $x$ values back to $-1$ (Eq. (2)), which always results in at least one Lyapunov exponent of $\infty$- ∞ [49]. Therefore, we perform a direct computation of the fractal dimension of the chaotic pseudo-attractor using the standard box-counting algorithm [54]. Using four-dimensional boxes of size $\epsilon=1/20,1/30,1/40$, we get that the box-counting dimension of the chaotic pseudo-attractor is $d\approx 1.84$ with an $R^{2}$ value of 0.9999, indicating that the chaotic pseudo-attractor is indeed fractal and strange. However, the pseudo-attractor’s dimension is still quite small for a geometric object embedded in four-dimensional space, which we attribute to the nonchaotic silence between bursts of spikes [see Fig. 1(b) and the left sides of Figs. 4(a) and 4(b)], as well as the aforementioned resetting mechanism.
(a) $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ vs. $k$
(b) $\lambda_{1}$ vs. $k$
Figure 5: Plots of the first three Lyapunov exponents at each point in the orbit of $\mathbf{X}_{0}=(-0.56,-3.25,-1,-3.25)$
We are also interested in the Lyapunov dimension
@@
and we assume the Lyapunov exponents are ordered from greatest to least [55]. The Kaplan-Yorke conjecture states that the Lyapunov dimension of typical attractors is equal to their fractal dimension [56]. For the chaotic pseudo-attractor orbit displayed in Fig. 4, if we naively calculate the Lyapunov dimension using the Lyapunov spectrum at one point in the orbit of an initial state that gets attracted to the chaotic pseudo-attractor, we will most likely not get something that coincides with the pseudo-attractor’s true fractal dimension. For example, if we consider the orbit generated from the initial state $\mathbf{X}_{0}=(-0.56,-3.25,-1,-3.25)$, the Lyapunov spectrum is approximately $\lambda\approx\{0.0165,-0.0105,-0.0810,-\infty\}$ at iteration $k=66000$, which is right before the state gets attracted to the nonchaotic spiking attractor [see Fig. 2(a)]. Using this, Eq. (12) indicates that the Lyapunov dimension of the chaotic pseudo-attractor is approximately $d_{l}\approx 2.07$, which is reasonably far from the pseudo-attractor’s true fractal dimension. The reason for this discrepancy lies in the fact that these chaotic dynamics are occurring on a pseudo-attractor, not a true attractor. Although the Kaplan-Yorke conjecture almost always holds for true attractors [57], the fact that we are dealing with a pseudo-attractor means that states do not spend enough time on the pseudo-attractor for their Lyapunov spectrums to capture its true geometry.
This fact can be easily seen in Fig. 5(a), where we plot the values of the first three Lyapunov exponents at all the iterations of the orbit of the initial state $\mathbf{X}_{0}=(-0.56,-3.25,-1,-3.25)$ that take place on the chaotic pseudo-attractor. Although the state evolves on the pseudo-attractor for tens of thousands of iterations, the figure shows that the Lyapunov exponents have not yet had time to converge to capture the true geometry of the pseudo-attractor, and we see many significant fluctuations in the $\lambda$ values, especially with $\lambda_{3}$. Of course, it is clear from Eqs. (12) and (13) that $\lambda_{3}$ doesn’t contribute the Lyapunov dimension, but Fig. 5(b) demonstrates that the maximal Lyapunov exponent $\lambda_{1}$ has also not converged; it is not oscillating around a mean value, but instead undergoing significant upwards and downwards trends. This explains this rare case where the fractal dimension and Lyapunov dimension of an attractor do not align. It also indicates that our computed Lyapunov exponents can at best be treated as a rough approximation for the true Lyapunov exponents of an orbit. Indeed, in the following sections, we use Lyapunov exponents mainly to determine whether an orbit is chaotic or nonchaotic, not as an exact quantifier for chaotic dynamics.
@@ IV Basin classification
We have established that a quasimultistability exists in this asymmetrically coupled neuron system, so we are now interested in analyzing the geometrical properties of the basins of attraction of the nonchaotic spiking attractor and strange pseudo-attractor. We begin by classifying the basins of attraction using the method of Sprott and Xiong [26], which we briefly outline here. Let $A=\{\mathbf{a}_{1},\mathbf{a}_{2},\ldots\}$ be an attractor  [^3] We assume $A$ is composed of an infinite number of points, but this method can be trivially altered for attractors of a finite number of points.. Define the mean or “center of mass” of $A$ to be
@@
(a) Maximal Lyapunov exponent color map
(b) Chaotic (white) and nonchaotic (black) basins
@@ IV.1 Two-dimensional basin slices
We now need a way to determine which attractor an initial state $\mathbf{X}_{0}\in S_{2}^{\prime}$ is initially attracted to. This is nontrivial because the nonchaotic spiking attractor and chaotic pseudo-attractor occupy similar regions of state space (see Figs. 3 and 4). Figure 2 illustrates that states attracted to the chaotic pseudo-attractor spend on the order of $10^{4}$ iterations in the pseudo-attractor before going to the spiking attractor. We find that testing other initial states yields similar results. Therefore, we use an indirect method of determining which basin an initial state is in, saying that an initial state that has a positive maximal Lyapunov exponent after 5,000 iterations is attracted to the chaotic pseudo-attractor, while an initial state that has a negative $\lambda_{1}$ after the same amount of time is attracted to the spiking attractor.
Using this method on a large number of initial states $\mathbf{X}_{0}\in S_{2}^{\prime}$, we present our maximal Lyapunov exponent results with a color map in Fig. 6(a). An immediate observation is the beautiful complexity of the different Lyapunov exponents in state space. This complexity is further emphasized when we color the basins properly in Fig. 6(b), with the basin of the chaotic pseudo-attractor (positive $\lambda_{1}$) shown in white and the basin of the nonchaotic attractor (negative $\lambda_{1}$) shown in black. Here, we can see that most of $S_{2}^{\prime}$ is taken up by the basin of the white chaotic pseudo-attractor, with some black regions, curves, and scattered points of stability scattered throughout. The most obvious of these features is the black line that goes across the diagonal of Fig. 6(b) from the bottom left to the top right. This line and its surrounding area are associated with initial states where the two individual neurons’ states are the same or very close to being the same. In this case, it makes sense that the neurons will immediately be synchronized, spiking together in the nonchaotic attractor.
@@
Because $S_{2}$ is two-dimensional, we also need to specify what is meant by the mean and standard deviation of the four-dimensional attractors within the two-dimensional plane $S_{2}$. Let us denote a generic four-dimensional attractor as $A_{4}$ and its associated basin as $\hat{A}_{4}$  [^4] In the asymmetrically coupled Rulkov neuron system, $A_{4}$ could denote either the nonchaotic spiking attractor or the chaotic pseudo-attractor.. From Eq. (14), if
@@
which is the projection of $\langle A_{4}\rangle$ onto the $(x_{1},x_{2})$ plane. Similarly, the true standard deviation of $A_{4}$ in four-dimensional state space is, from Eq. (15),
@@
Following Eq. (16), for our classification of the basin slices $\hat{A}_{4}\cap S_{2}$, we will consider the normalized two-dimensional distance $\xi_{2}$ from $\langle A_{2}\rangle$ of some $\mathbf{X}=(x_{1},-3.25,x_{2},-3.25)\in S_{2}$ to be,
@@
To classify the basin slices of the chaotic pseudo-attractor and nonchaotic spiking attractor, we want to find the two probability functions $P_{w}(\xi_{2})$ and $P_{b}(\xi_{2})$ associated with the white (chaotic) and black (nonchaotic) basins, respectively, which we accomplish using an iterative Monte Carlo method. Specifically, we first randomly select states from the open disk of radius $\xi_{2}=1$ centered at $\langle A_{2}\rangle$ and calculate the maximal Lyapunov exponents of the initial states’ orbits to determine which basin they end up in. Using this, we calculate the value of $P(1)$. Then, starting from $k=0$, we repeat the same process for the two-dimensional shell centered at $\langle A_{2}\rangle$ with inner radius $\xi_{2}=2^{k}$ and outer radius $\xi_{2}=2^{k+1}$ to calculate the value of $\Delta P(2^{k})$. Then, we use Eq. (21) to find $P(2^{k+1})$ and repeat the process for $k+1$. One important numerical note is that, given we are testing values of $\xi_{2}=2^{k}$, the maximum value of $k$ must be chosen with care. On the one hand, since we are interested in the limit $\xi\to\infty$, $k$ must be large enough that we capture the geometry of the basins far away from the attractors. On the other hand, since we are only iterating 5000 times, it also must be small enough that initial states go to an attractor fast enough to get an accurate Lyapunov exponent calculation on it. With a computer experiment, we determine that a good maximum value of $k$ is 11.
@@
Table 1: Some approximate $P(\xi_{2})$ values of the asymmetrically coupled Rulkov neuron system’s chaotic (white) and nonchaotic (black) basins
In Table 1, we display our numerical results for the values of $P_{w}(\xi_{2})$ and $P_{b}(\xi_{2})$, where we neglect values of $k$ from 0 to 5 since we are interested in large values of $\xi_{2}$. We can immediately see that $P_{w}(\xi_{2})$ is behaving a bit oddly, as Eq. (20) indicates that $P(\xi)$ functions should either decrease or stay the same. For this reason, we first turn our attention to $P_{b}(\xi_{2})$, which is behaving as expected. Performing a linear regression calculation on $\log_{2}P_{b}(\xi_{2})$ against $\log_{2}\xi_{2}$, we get that
@@
Since $\gamma\approx 0.2957$ is between 0 and $n=2$, we can conclude that the basin slice of the nonchaotic spiking attractor is Class 3, meaning it extends to infinity in some directions but takes up increasingly small fractions of state space as we move outwards. This makes sense because the farther away an initial state is from the nonchaotic spiking attractor, the lower the chance there will be for the neurons to immediately synchronize with each other. In the limit $\xi_{2}\to\infty$, we expect that the only initial conditions that get attracted to the spiking attractor immediately are the ones where $x_{1,0}=x_{2,0}$, which is the diagonal in Fig. 6. For this reason, we can conclude that the black basin slice in $S_{2}$ has finite measure since the line $x_{1,0}=x_{2,0}$ doesn’t contribute to the measure of the two-dimensional black basin slice.
@@
by the definitions of the spiking attractor and chaotic pseudo-attractor: if an initial state does not immediately fall into a nonchaotic spiking orbit with negative $\lambda_{1}$, then it goes to a chaotic orbit with positive $\lambda_{1}$. This fact can be confirmed by adding the data horizontally across Table 1, which gives values very close to 1. Therefore, even though $P_{w}(2^{11})$ is less than $P_{w}(2^{10})$ due to the aforementioned difficulty of calculating Lyapunov exponents for orbits of initial states far away from their attractors, the values of $P_{w}(\xi_{2})$ are indeed approaching 1. This can be seen by exponentiating both sides of Eq. (30):
@@ IV.2 Four-dimensional basins
using the definitions from Eqs. (25) and (27). This is similar to the process for the two-dimensional slices, with a few noteworthy changes. One nontrivial aspect of this is choosing a random state $\mathbf{X}_{0}$ from a four-dimensional open ball with radius $\xi_{4}$ centered at $\langle A_{4}\rangle$. In two-dimensional space with an open disk $|\mathbf{x}-\langle A_{2}\rangle|<\xi_{2}$, we simply choose a random $r\in[0,\xi_{2})$ and a random $\phi\in[0,2\pi)$, then say the random initial state is $\mathbf{x}_{0}=\langle A_{2}\rangle+(r\cos\phi,r\sin\phi)$. In four-dimensional space, we can use the analog to spherical coordinates in four-dimensions, where four coordinates $r\in[0,\xi_{4})$, $\theta_{1}\in[0,\pi)$, $\theta_{2}\in[0,\pi)$, and $\phi\in[0,2\pi)$ are chosen randomly. Then, a random initial state $\mathbf{X}_{0}\in S(\xi_{4})$ is given by [60]
@@
Table 2: Some approximate $P(\xi_{4})$ values of asymmetrically coupled Rulkov neuron system’s chaotic (white) and nonchaotic (black) basins
@@
(b) Chaotic (white) and nonchaotic (black) basins
@@
The results for the probability function values of the white basin $P_{w}(\xi_{4})$ are shown in Table 2. In the table, the $P_{b}(\xi_{4})$ values are indirectly calculated using $P_{b}(\xi_{4})=1-P_{w}(\xi_{4})$, which follows from Eq. (31). Observing the values in Table 2, it appears that $P_{w}(\xi_{4})$ and $P_{b}(\xi_{4})$ remain relatively constant as $\xi_{4}$ is varied, which is a rather unexpected result given our analysis of the two-dimensional basin slices. Running linear regressions on the log-log plots of both $P_{w}(\xi_{4})$ and $P_{b}(\xi_{4})$, we find that
@@
with $R^{2}=0.386$. These low $R^{2}$ values confirm that there is no upward or downward trend in $P_{w}(\xi_{4})$ or $P_{b}(\xi_{4})$. This means that, within the numerical variability of the Monte Carlo method, the value of the basin classification parameter $\gamma$ is effectively 0 for both the white and black basins. From Eqs. (35) and (36), the value of the other basin classification parameter $P_{0}$ is $P_{0}=0.8517$ for the white basin and $P_{0}=0.1482$ for the black basin. Both of these are between 0 and 1, so both the white and black basins are Class 2, meaning they both occupy a fixed fraction of four-dimensional state space. This makes sense considering the data in Table 2, but again, it is an unexpected result.
@@
Our explanation for why the white and black basins are Class 2 relies on the fact that the nonchaotic Rulkov map is a slow-fast system. Because $x$ is a fast variable and $y$ is a slow variable, $x_{1}$ and $x_{2}$ adjust much faster than $y_{1}$ and $y_{2}$, so if we consider some initial state $\mathbf{X}_{0}=(x_{1,0},y_{1,0},x_{2,0},y_{2,0})$ some distance $\xi_{4}$ away from either attractor $A_{4}$, $y_{1}$ and $y_{2}$ will slowly drift towards $\langle y_{1}\rangle$ and $\langle y_{2}\rangle$, respectively, and $x_{1}$ and $x_{2}$ will easily keep up with their slow attraction. By the time $y_{1}$ and $y_{2}$ reach the region around $\langle y_{1}\rangle$ and $\langle y_{2}\rangle$, $x_{1}$ and $x_{2}$ will also be near $\langle x_{1}\rangle$ and $\langle x_{2}\rangle$, respectively. This will occur regardless of how much farther away $x_{1,0}$ and $x_{2,0}$ are from $\langle x_{1}\rangle$ and $\langle x_{2}\rangle$ compared to the distance from $y_{1,0}$ and $y_{2,0}$ to $\langle y_{1}\rangle$ and $\langle y_{2}\rangle$, respectively, since the fast variables $x_{1}$ and $x_{2}$ evolve so much more quickly than the slow variables $y_{1}$ and $y_{2}$. Thus, once $y_{1}$ and $y_{2}$ get close to the attractors, we are effectively starting in $S_{2}^{\prime}$ (or a square of similar size close and parallel/skew to it). Indeed, comparing the fractions of state space that the white ($P_{0}=0.8517$) and black ($P_{0}=0.1482$) basins take up to the values of $P_{w}(\xi_{2})$ and $P_{b}(\xi_{2})$ for the small values of $\xi_{2}$ in Table 1, we find that these align quite well.
@@
Although we cannot visualize the entirety of these four-dimensional basins, we can get a grasp of this Class 2 basin behavior by graphing a different two-dimensional slice of the basins, namely, one where $y_{1,0}$ and $y_{2,0}$ are varied but $x_{1,0}$ and $x_{2,0}$ are fixed. In Fig. 7, we graph a large range of initial $y$ values $-256\leq y_{1,0}\leq 256$ and $-256\leq y_{2,0}\leq 256$ while fixing the initial $x$ values at $x_{1,0}=-1$ and $x_{2,0}=1$. In Fig. 7(b), we can see a dispersed, seemingly random distribution of white and black basin points with no defined pattern in their arrangement besides some subtle outward extending darker and lighter regions. Similarly, in Fig. 7(a), the maximal Lyapunov exponent also appears random. This apparent randomness makes sense from our discussion of how far-away points evolve toward the attractors. Because of how slow $y_{1}$ and $y_{2}$ move towards their eventual values, by the time the state reaches a square close to $S_{2}^{\prime}$, the fast variables $x_{1}$ and $x_{2}$ have evolved so much that they are essentially randomized. In other words, any difference in $y_{1}$ and $y_{2}$ between two initial states will result in substantially different $x_{1}$ and $x_{2}$ values once the state reaches a square close to $S_{2}^{\prime}$. This essentially turns the initial state from $S_{4}$ into a random initial state in $S_{2}^{\prime}$ (see Fig. 6).
The important observation from Fig. 7(b) is that despite the angular variances in the distribution of white and black points, the distribution appears to remain the same regardless of the radial distance from the center. Even though the plot spans a large range of different $y$ values, this consistent distribution characteristic of Class 2 basins is present: the fraction of state space Class 2 basins take up doesn’t depend on the distance from their attractors. Figure 7(b) also lacks the black diagonal that appears in Fig. 6(b). This line of stability doesn’t appear in this graph because, as previously mentioned, it is defined by